# Phase space analysis of CCDM cosmologies

### v1

@@ III Phase space Analysis
In order to proceed with the dynamical analysis, we introduce the following dimensionless variables
$x=\frac{\kappa^{2}\rho_{\rm dm}}{3H^{2}},\quad\mbox{and}\quad z=\frac{H_{0}}{H_{0}+H},$ (24)
where $H_{0}$ represents the present value of the Hubble rate. [^3] Note that without any loss of generality, one can use $H_{*}$ (the value of the Hubble parameter at $t=t_{*}$) instead of $H_{0}$ and this will not change the physics. Now, using the dimensionless variables in (24), the gravitational equations for one fluid or two-fluid systems as described in section II can be converted into the autonomous systems. Considering the general matter creation rate $\Gamma$ in (14), the decelerating parameter, $q\equiv-1-\dot{H}/H^{2}$, for the one-fluid system reads
$q=\frac{1}{2}\left(1-\frac{\Gamma}{H}\right),$ (25)
and for two-fluid system in terms of the dimensionless variables in (24), takes the form
$q=\frac{1}{2}\left[1+3w(1-x)-\frac{\Gamma}{H}x\right].$ (26)
@@
III.1 Model: $\Gamma=\Gamma_{0}$
In this section we shall discuss the effects of the constant particle creation rate $\Gamma=\Gamma_{0}$ considering a one fluid system and a two-fluid system.
III.1.1 One fluid system
From the first Friedmann eqn. (7), it is clear that the DM density parameter, $\Omega_{\rm dm}=\kappa^{2}\rho_{\rm dm}/3H^{2}$ always takes the value $1$. Thus, we are interested on the evolution of the $z$ variable. The dynamics of the matter creation model for $\Gamma=\Gamma_{0}$ can be described by the following one dimensional equation
$z^{\prime}=\frac{3}{2}z\left(1-z-\frac{\alpha z}{3}\right),$ (27)
where $\alpha=\Gamma_{0}/H_{0}>0$ is the dimensionless parameter and the prime denotes the derivative with respect to $N=\ln(a/a_{0})$. The decelerating parameter $q$ adopts the form $q=\frac{1}{2}\left(1-\frac{\alpha z}{1-z}\right)$ which at present time (i.e. at $z=\frac{1}{2}$) reduces to the form $q=\frac{1-\alpha}{2}$. Thus, at present time in order to realize an accelerating phase, the required condition is $\alpha>1$. The relevant critical points of the dynamical system (27) and their cosmological features are summarized in Table 1. In this case, we obtain only two critical points, namely, $A^{\prime}_{0},~{}A^{\prime}_{1}$ in which the first component corresponds to DM density parameter and second component represents the $z$ variable. Here $A^{\prime}_{0}$ is a DM dominated unstable point showing the decelerating phase, while $A^{\prime}_{1}$ is also a DM dominated stable critical point giving accelerating phase. The directions of the vector field involving in the system (27) are presented in the upper left plot of the Fig. 1 showing a transition of our universe from its past decelerating phase to the current accelerating one.
III.1.2 Two-fluid system
For the two-fluid system, considering the dimensionless variables, $x$, $z$ defined in eqn. (24), the constraint from the Friedmann equation (11) is, $\Omega_{\rm dm}+\Omega_{f}=1$ where $\Omega_{f}=\kappa^{2}\rho_{f}/(3H^{2})$. Now, using these dimensionless variables, one can write down the autonomous system as follows
$x^{\prime}=$ $x(1-x)\left(3w+\frac{\alpha z}{1-z}\right),$ (28a)
where $\alpha=\Gamma_{0}/H_{0}$ is a parameter which takes only positive values. For this model our physical region is a unit square and we define it as, $R_{1}=[0,1]^{2}$. The above autonomous system has a singularity at $z=1$, but we want access the phase space analysis of the points which correspond to $z=1$. So, we multiply the right hand sides of (28a), (28b) by the non-negative factor $(1-z)$ and regularize the autonomous system. Therefore, the reduced autonomous system can be read as:
$x^{\prime}=$ $x(1-x)\left(3w(1-z)+\alpha z\right),$ (29a)
$z^{\prime}=$ $\frac{3}{2}z(1-z)\bigg{[}\bigg{(}1+w(1-x)\bigg{)}(1-z)-\frac{\alpha xz}{3}\bigg{]},$ (29b)
From the above autonomous system, one can easily conclude that $x=0,~{}x=1,~{}z=0$ and $z=1$ are invariant manifolds which indicate that our physical region is positively invariant, i.e. if we take any orbit from $R_{1}$, it never leaves the domain. Here, the decelerating parameter defined in (26) is simplified to $q=\frac{1}{2}\left[1+3w(1-x)-\frac{\alpha xz}{1-z}\right]$. From this relation, one can quickly derive that in order to obtain the accelerating expansion at present time, i.e. at $z=1/2$, the viable orbits must satisfy the relation $(1-\alpha)x+(1+3w)(1-x)<0$.
Now, in order to understand the phase space dynamics of the autonomous system (29), we find the critical points and investigate their stability conditions. The critical points and their characteristics are summarized in Table 1. In this case, we find five discrete critical points ($A_{0}$, $A_{1}$, $A_{2}$, $A_{3}$, and $A_{4}$) and three critical lines ($A_{5}$, $A_{6}$ and $A_{7}$). Since the equation of state of the second fluid, $w$, is involved and as $w$ can take any value (either non-negative or negative), therefore, in order to understand how the nature of the second fluid affects the architecture of the phase space dynamics, we have paid special attention to the cases with $w\geq 0$ indicating the normal fluid and $w<0$ representing the hypothetical fluid. In the following we describe the qualitative features of the above dynamical system for different values of $w$.
One fluid system
Critical point Existence Stability $\Omega_{\rm dm}$ $q$ Acceleration
$A^{\prime}_{0}(1,0)$ Always Always Unstable 1 $\frac{1}{2}$ No
$A^{\prime}_{1}\left(1,\frac{3}{3+\alpha}\right)$ $\alpha>0$ Always Stable 1 $-1$ Yes
Two-fluid system
Critical point Existence Eigenvalue Stability $\Omega_{f}$ $\Omega_{\rm dm}$ $q$ Acceleration
$A_{0}(0,0)$ Always $\left(3w,\frac{3}{2}(1+w)\right)$ Stable if $w<-1$; 1 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Saddle if $-1<w<0$;
@@ III.1.2 Two-fluid system
$A_{2}(0,1)$ Always $(\alpha,0)$ Non-hyperbolic Saddle if $w>-1$ 1 0 Undefined Undetermined
Unstable if $w<-1$
$A_{3}(1,1)$ Always $\left(-\alpha,\frac{\alpha}{2}\right)$ Always Saddle 0 1 $\infty$- ∞ Yes
$A_{4}\left(1,\frac{3}{3+\alpha}\right)$ $\alpha>0$ $\left(-\frac{3\alpha(1+w)}{3+\alpha},-\frac{3\alpha}{2(3+\alpha)}\right)$ Stable if $w>-1$ 0 1 $-1$ Yes
Saddle if $w<-1$
$A_{5}\left(0,z_{c}\right)$ $w=-1$ $\left((3+\alpha)z_{c}-3,0\right)$ Stable if $z_{c}<\frac{3}{3+\alpha}$; $1$ 00 $-1$ Yes
@@
$A_{6}\left(x_{c},\frac{3}{3+\alpha}\right)$ $w=-1$ $\left(0,-\frac{3\alpha x_{c}}{2(3+\alpha)}\right)$ Stable $1-x_{c}$ $x_{c}$ $-1$ Yes
@@
Table 1: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (27) and (29) with the matter creation rate $\Gamma=\Gamma_{0}$.
@@
Figure 1: Description of the phase space of the matter creation scenarios (one fluid and two-fluid systems) in which the matter creation rate is constant, i.e. $\Gamma=\Gamma_{0}$. Upper Left Plot: This is the vector plot of the one dimensional system (27). Here, we adopt $\alpha=3$ but for other values of $\alpha>0$ similar type of plot can be found . Upper Right Plot: The phase plot of the system (29) when we have assumed $w=1.5$ and $\alpha=2$. For other values of $w>0$ and $\alpha>0$, we can also obtain similar phase space structure. Lower Left Plot: The phase space of the system (29) when the equation of state $w$ takes the value 00. Here we use $\alpha=2$ but any positive value of $\alpha$ gives same type of phase portrait. Lower Right Plot: The phase plot of the system (29) considering $w=-0.8$ and $\alpha=3$. Also, we can get similar type of graphics for any positive value of $\alpha$ and negative value of $w$ in the interval $(-1,0)$. In all three two-dimensional plots, green region corresponds to the decelerating phase ($q>0$), cyan region represents the accelerating phase with $-1<q<0$ and the yellow region corresponds to the super accelerating phase (i.e. $q<-1$). The black dotted curve separating the cyan and yellow regions corresponds to $q=-1$.
Figure 2: This figure depicts the phase space of the two-fluid system for the constant matter creation rate $\Gamma=\Gamma_{0}$ but here we consider the case with $w\leq-1$. Left Plot: The phase space of the system (29) when the equation of state $w$ takes the value $-1$. Here we use $\alpha=3$ but any positive value of $\alpha$ gives similar type of phase portrait. Right Plot: The phase plot of the system (29) when we assume $w=-1.5$ and $\alpha=3$. For other values of $w<-1$ and $\alpha>0$, we can also obtain similar graphics. In both the plots, green region corresponds to the decelerating phase ($q>0$), cyan region represents the accelerating phase with $-1<q<0$ and the yellow region corresponds to the super accelerating phase (i.e. $q<-1$). The black dotted curve separating the cyan and yellow regions (in the left plot, this separating curve is not visible because of the line of critical points) corresponds to $q=-1$.
  1. 1.
@@
    When $w>0$: In this case, we obtain five critical points, namely, $A_{0}$, $A_{1}$, $A_{2}$, $A_{3}$ and $A_{4}$ which always exist in our region $R_{1}$. Eigenvalues evaluated at $A_{0}$ are $3w$, $\frac{3}{2}(1+w)$ which are positive, hence, this indicates that $A_{0}$ is an unstable critical point dominated by the normal fluid and it represents a decelerating phase of the universe. At $A_{1}$, the eigenvalues are $-3w$, $\frac{3}{2}$, that means one eigenvalue is positive while the other is negative (since $w>0$). So, the critical point $A_{1}$ is saddle in nature which corresponds to the DM dominated decelerating phase. The critical point $A_{2}$ has one eigenvalue 00. So, we cannot say about stability by applying the linear stability theory and that is why we have applied the center manifold theory. The center manifold is $x=0$ which coincides with the eigenvector ($z$ axis) corresponding to a zero eigenvalue. Clearly, $z^{\prime}$ is positive along $z$ axis which gives $A_{2}$ another saddle point. As the decelerating parameter is undefined at $A_{2}$, hence, we are unable to predict whether this represents an accelerating phase or the reverse. Since, $A_{3}$ has two eigenvalues $-\alpha$, $\frac{\alpha}{2}$ and they are of opposite signs, we can conclude that it is a saddle point. This critical point is a DM dominated accelerated solution. The most interesting critical point is $A_{4}$ which has $\Omega_{dm}=1$ and $q=-1$, i.e. DM dominated accelerating solution. As both the eigenvalues are negative, the critical point $A_{4}$ is the only globally stable point in our domain which is shown in the upper right plot of Fig. 1 where all orbits leave the decelerating phase, then enter into an accelerating phase and finally end in a DM dominated accelerating phase $(q=-1)$. Besides these features we observe that a set of orbits crosses the $q=-1$ curve, enter into a super accelerating phase $(q<-1)$ and finally finish with $q=-1$ phase, which indicates the slowing down of the cosmic acceleration.
  2. 2.
@@
    When $w=0$: In this scenario besides the three isolated critical points $A_{2},~{}A_{3}$ and $A_{4}$ given in Table 1, we find one critical line $A_{7}(x_{c},0)$ for which the critical points $A_{0}$, $A_{1}$ lie in this critical line. In the phase space $R_{1}$, one can obtain $z^{\prime}>0$ and $z^{\prime}<0$ for $z<\frac{3}{3+\alpha x}$ and for $z>\frac{3}{3+\alpha x}$ respectively, and $x^{\prime}$ takes always positive value (since $\alpha>0$). By the above argument, one can conclude that the critical line $A_{7}$ shows unstable nature, $A_{1},~{}A_{2}$ behave like saddle points and $A_{3}$ is the only globally stable point. Here, $A_{3},~{}A_{4}$ correspond to accelerating phase and $A_{7}$ gives decelerating phase, while decelerating parameter is undefined at $A_{2}$. The phase space diagram is highlighted in the lower left plot of the Fig. 1 where one can notice that the universe exits from its past decelerating phase and enters into the DM dominated accelerating phase. Along with these results, a set of orbits traverses the decelerating phase $(q>0)$, then enters into an accelerating phase with $-1<q<0$ but successively these orbits enter into a super accelerating phase $(q<-1)$, and finally the acceleration slows down and it continues with $q=-1$.
  3. 3.
@@
    When $-\frac{1}{3}\leq w<0$: In this interval of $w$, the dynamical system (29) gives five critical points, namely, $A_{0},~{}A_{1},~{}A_{2},~{}A_{3}$ and $A_{4}$. Now we can conclude about the stability of the critical points via investigation of the sign of the eigenvalues and the direction of the flow on the boundary of the unit square. It is clear that $A_{0},~{}A_{2},~{}A_{3}$ are saddle by behavior and $A_{1}$ is unstable. Here, $A_{4}$ is the only globally stable point in the phase space. The critical points $A_{3},~{}A_{4}$ correspond to matter dominated accelerating phases, while $A_{1}$ represents the matter dominated decelerating phase. Note that, $A_{0},~{}A_{2}$ both are second fluid dominated critical points but again $A_{0}$ shows deceleration and we can not say about the acceleration at $A_{2}$. The phase space trajectories are similar to the lower right plot of the Fig. 1. In this case, the universe quits its past DM dominated decelerating phase and closes in accelerating DM dominated phase. Again, some special orbits tracing accelerating phase $(-1<q<0)$ and super accelerating phase $(q<-1)$, finish in the phase where $q=-1$. The slowing down of cosmic acceleration is also realized in this case.
  4. 4.
@@
    When $-1<w<-\frac{1}{3}$: This case is similar to the above case where $w$ lies in the interval $\left[-\frac{1}{3},0\right)$. The only difference with the above case is in the decelerating parameter where in the above case we get $q\geq 0$ at $A_{0}$ but here, $q<0$ at $A_{0}$ which represents acceleration. The lower right plot of Fig. 1 exhibits the qualitative features of the critical points.
  5. 5.
@@
    When $w=-1$: In this case, we have two critical lines $A_{5}(0,z_{c})$ and $A_{6}\left(x_{c},\frac{3}{3+\alpha}\right)$ along with the critical points $A_{0},~{}A_{1},~{}A_{2},~{}A_{3},~{}A_{4}$ depicted in the Table 1. Here, $A_{0},~{}A_{2}$ belong to the critical line $A_{5}$ and $A_{6}$ contains the point $A_{4}$. In the phase space above the $z=3/(3+\alpha)$ line, $x^{\prime}$, $z^{\prime}$ are positive and negative respectively, and below the line $z=3/(3+\alpha)$, $x$ and $z$ are decreasing and increasing respectively. Again, the separatrix $z=\frac{3}{\alpha}(1-z)\sqrt{1-x}$ connecting the critical points $A_{1}$ and $\left(0,3/(3+\alpha)\right)$ divide the phase space below the line $z=3/(3+\alpha)$ into two regions: the orbits below the separatrix converge to the part of the critical line $A_{5}$ where $z<\frac{3}{3+\alpha}$ and the orbits above the separatrix approach the critical line $A_{6}$. As a result, $A_{6}$ behaves like an accelerating late-time attractor where DM and DE both may co-exist. So, the critical point $A_{6}$ has the potentiality to alleviate the coincidence problem. Again, the critical line $A_{5}$ where $z>\frac{3}{3+\alpha}$ are unstable in nature but the part of the critical line $A_{5}$ with $z<\frac{3}{3+\alpha}$ represent late time accelerating attractor dominated by the DE only. The point $A_{3}$ is saddle which shows matter dominated accelerated phase, while $A_{1}$ corresponds to matter dominated decelerating unstable point. The qualitative behavior are highlighted in the left plot of Fig. 2. Here, we have two possibilities: one is, above the separatrix, our universe shows the transition from its past DM dominated decelerating phase to the accelerating scaling solutions and on the other hand, below the separatrix, the universe follows DM dominated decelerating phase to completely DE dominated accelerating phase.
  6. 6.

### v2

@@ III Phase space Analysis
In this section we present a detailed phase space analysis of the two-fluid matter creation scenarios considering various matter creation rates. As far as we are concerned with the literature, no systematic analyses of such matter creation models are present, and hence, such analyses and results are completely new in this direction. In order to proceed with the dynamical analysis, we introduce the following dimensionless variables
$x=\frac{\kappa^{2}\rho_{\rm dm}^{\rm eff}}{3H^{2}},\quad\mbox{and}\quad z=\frac{H_{0}}{H_{0}+H},$ (20)
where $H_{0}$ represents the present value of the Hubble rate. [^4] Note that without any loss of generality, one can use $H_{*}$ (the value of the Hubble parameter at $t=t_{*}$) instead of $H_{0}$ and this will not change the physics. Therefore, the variables $x$ and $z$, taking values in the intervals $0\leq x\leq 1$ and $0\leq z\leq 1$, define the phase space as a unit square ${\bf R}=[0,1]^{2}$. Now, using the dimensionless variables in (20), the gravitational equations describing the dynamics of the fluids can be converted into the autonomous systems. For any matter creation rate $\Gamma$, the decelerating parameter, $q\equiv-1-\dot{H}/H^{2}$, in terms of the dimensionless variables in (20), takes the form
$q=\frac{1}{2}\left[1+3w(1-x)-\frac{\Gamma xz}{(1-z)H_{0}}\right].$ (21)
@@
Critical point Existence Eigenvalue Stability $\mathbf{\Omega_{f}}$ $\mathbf{\Omega_{\rm dm}^{\rm eff}}$ $\mathbf{q}$ Acceleration
$A_{0}(0,0)$ Always $\left(3w,\frac{3}{2}(1+w)\right)$ Stable if w<1w<-1; 1 0 $\frac{1}{2}(1+3w)$ $w<-\frac{1}{3}$
Saddle if 1<w<0-1<w<0;
@@
$A_{2}(0,1)$ Always $(\alpha,0)$ Non-hyperbolic Saddle if w>1w>-1 1 0 Undefined Undetermined
Unstable if w<1w<-1
$A_{3}(1,1)$ Always $\left(-\alpha,\frac{\alpha}{2}\right)$ Always Saddle 0 1 -$\infty$ Yes
$A_{4}\left(1,\frac{3}{3+\alpha}\right)$ $\alpha>0$ $\left(-\frac{3\alpha(1+w)}{3+\alpha},-\frac{3\alpha}{2(3+\alpha)}\right)$ Stable if w>1w>-1 0 1 1-1 Yes
Saddle if w<1w<-1
$A_{5}\left(0,z_{c}\right)$ w=1w=-1 $\left((3+\alpha)z_{c}-3,0\right)$ Stable if $z_{c}<\frac{3}{3+\alpha}$; 11 0 1-1 Yes
@@
$A_{6}\left(x_{c},\frac{3}{3+\alpha}\right)$ w=1w=-1 $\left(0,-\frac{3\alpha x_{c}}{2(3+\alpha)}\right)$ Stable $1-x_{c}$ $x_{c}$ 1-1 Yes
@@
Table 1: Summary of the critical points, their existence, stability and the values of the cosmological parameters at those points for the dynamical system (23) with the matter creation rate $\Gamma=\Gamma_{0}$.
III.1 Model: $\Gamma=\Gamma_{0}$
In this section we shall discuss the effects of the constant particle creation rate $\Gamma=\Gamma_{0}$ considering two-fluid system.
For the two-fluid system, considering the dimensionless variables, $x$, $z$ defined in eqn. (20), the constraint from the Friedmann equation (7) with $k=0$ is, $\Omega_{\rm dm}^{\rm eff}+\Omega_{f}=1$ where $\Omega_{f}=\kappa^{2}\rho_{f}/(3H^{2})$ and $\Omega_{\rm dm}^{\rm eff}=\kappa^{2}\rho_{\rm dm}^{\rm eff}/(3H^{2})$. [^5] Note that $\Omega_{f0}$, $\Omega_{\rm dm,0}^{\rm eff}$ are the present day values of the density parameters. Now, using these dimensionless variables, one can write down the autonomous system as follows
$x^{\prime}=$ $x(1-x)\left(3w+\frac{\alpha z}{1-z}\right),$ (22a)
where prime denotes the derivative with respect to $N=\ln(a)$ [^6] From now on, prime attached to any variable will refer to the derivative with respect to $N$. and $\alpha=\Gamma_{0}/H_{0}$ is a parameter which takes only positive values. To address the singularity at $z=1$ of the dynamical system (22), we introduce a new time variable $\tau$, defined as $dN=(1-z)d\tau$. Consequently, the reduced autonomous system which is topologically equivalent [178, 179, 180, 181, 182, 183, 184, 185] to the autonomous system (22) can be read as:
$\frac{dx}{d\tau}=$ $x(1-x)\left(3w(1-z)+\alpha z\right),$ (23a)
$\frac{dz}{d\tau}=$ $\frac{3}{2}z(1-z)\bigg[\bigg(1+w(1-x)\bigg)(1-z)-\frac{\alpha xz}{3}\bigg],$ (23b)
From the above autonomous system, one can easily conclude that x=0,x=1,z=0x=0,~x=1,~z=0 and $z=1$ are invariant manifolds which indicate that our physical region is positively invariant, i.e. if we take any orbit from $\bf R$}, it never leaves the domain. Here, the decelerating parameter defined in (21) is simplified to $q=\frac{1}{2}\left[1+3w(1-x)-\frac{\alpha xz}{1-z}\right]$. From this relation, one can quickly derive that in order to obtain the accelerating expansion at present time, i.e. at $z=1/2$, the viable orbits must satisfy the relation $(1-\alpha)x+(1+3w)(1-x)<0$.
Now, in order to understand the phase space dynamics of the autonomous system (23), we find the critical points and investigate their stability conditions. The critical points and their characteristics are summarized in Table 1. In this case, we find five discrete critical points ($A_{0}$, $A_{1}$, $A_{2}$, $A_{3}$, and $A_{4}$) and three critical lines ($A_{5}$, $A_{6}$ and $A_{7}$). Since the EoS of the second fluid, $w$, is involved and as $w$ can take any value (either non-negative or negative), therefore, in order to understand how the nature of the second fluid affects the architecture of the phase space dynamics, we have paid special attention to the cases with $w\geq 0$ indicating the normal fluid and $w<0$ representing the hypothetical fluid. In the following we describe the qualitative features of the above dynamical system for different values of $w$.
@@ III.1 Model: $\Gamma=\Gamma_{0}$
Figure 1: Description of the phase space of the matter creation scenarios in which the matter creation rate is constant, i.e. $\Gamma=\Gamma_{0}$. Upper Left Plot: The phase plot of the system (23) when we have assumed $w=0.1$ and $\alpha=2$. For other values of $w>0$ and $\alpha>0$, we can also obtain similar phase space structure. Upper Middle Plot: The phase space of the system (23) when the EoS $w$ takes the value 0. Here we use $\alpha=2$ but any positive value of $\alpha$ gives same type of phase portrait. Upper Right Plot: The phase plot of the system (23) considering w=0.8w=-0.8 and $\alpha=3$. Also, we can get similar type of graphics for any positive value of $\alpha$ and negative value of $w$ in the interval (1,0)(-1,0). Lower Left Plot: The phase space of the system (23) when the EoS $w$ takes the value 1-1. Here we use $\alpha=3$ but any positive value of $\alpha$ gives similar type of phase portrait. Lower Right Plot: The phase plot of the system (23) when we assume w=1.5w=-1.5 and $\alpha=3$. For other values of w<1w<-1 and $\alpha>0$, we can also obtain similar graphics. In all five two-dimensional plots, green region corresponds to the decelerating phase ($q>0$), cyan region represents the accelerating phase with 1<q<0-1<q<0 and the yellow region corresponds to the super accelerating phase (i.e. q<1q<-1). The black dotted curve separating the cyan and yellow regions (in the left plot, this separating curve is not visible because of the line of critical points) corresponds to q=1q=-1.
Figure 2: We provide the evolution plots for the dynamical system (23), showing the behavior of the DM density parameter $\Omega_{\rm dm}^{\rm eff}$, the second fluid density parameter $\Omega_{f}$, the deceleration parameter $q$, and the effective equation of state of DM $w_{\rm dm}^{\rm eff}$ under the constant particle creation rate $\Gamma=\Gamma_{0}$. Upper Left Plot: This figure is drawn with the model parameters $w=0.1$, $\alpha=1$, and it shows that the final fate of the universe is late time accelerated evolution in cosmological constant era dominated by DM. Upper Right Plot: This figure is produced for the parameters values $w=0$ and $\alpha=1$. Here also, the final fate of the universe is late time accelerated evolution in cosmological constant era dominated by DM. Lower Left Plot: We have used w=1w=-1 and $\alpha=1$. This plot depicts that the ultimate fate of the universe is found to be late time accelerated evolution in cosmological constant epoch which is achieved after the occurrence of a decelerated DM dominated phase. Lower Right Plot: We have taken w=1.3w=-1.3 and $\alpha=1$. This plot highlights that the ultimate fate of the universe is found to be late time DE dominated accelerated evolution in phantom epoch which is achieved after the occurrence of a decelerated DM dominated phase.
[…]
@@
When $w>0$, we obtain five critical points, namely, $A_{0}$, $A_{1}$, $A_{2}$, $A_{3}$ and $A_{4}$ which always exist in our region $R$. Eigenvalues evaluated at $A_{0}$ are $3w$, $\frac{3}{2}(1+w)$ which are positive, hence, this indicates that $A_{0}$ is an unstable critical point dominated by the normal fluid and it represents a decelerating phase of the universe. At $A_{1}$, the eigenvalues are 3w-3w, $\frac{3}{2}$, that means one eigenvalue is positive while the other is negative (since $w>0$). So, the critical point $A_{1}$ is saddle in nature which corresponds to the DM dominated decelerating phase. The critical point $A_{2}$ has one eigenvalue 0. So, we cannot say about stability by applying the linear stability theory and that is why we have applied the center manifold theory. The center manifold is $x=0$ which coincides with the eigenvector ($z$ axis) corresponding to a zero eigenvalue. Clearly, $z^{\prime}$ is positive along $z$ axis which gives $A_{2}$ another saddle point. As the decelerating parameter is undefined at $A_{2}$, hence, we are unable to predict whether this represents an accelerating phase or the reverse. Since, $A_{3}$ has two eigenvalues $\alpha$, $\frac{\alpha}{2}$ and they are of opposite signs, we can conclude that it is a saddle point. This critical point is a DM dominated accelerated solution. The most interesting critical point is $A_{4}$ which has $\Omega_{dm}^{\rm eff}=1$ and q=1q=-1, i.e. DM dominated accelerating solution. As both the eigenvalues are negative, the critical point $A_{4}$ is the only globally stable point in our domain which is shown in the upper left plot of Fig. 1 where all orbits leave the decelerating phase, then enter into an accelerating phase and finally end in a DM dominated accelerating phase (q=1)(q=-1), as evidenced in the upper left plot in Fig. 2. Besides these features we observe that a set of orbits crosses the q=1q=-1 curve, enter into a super accelerating phase (q<1)(q<-1) and finally finish with q=1q=-1 phase, which indicates the slowing down of the cosmic acceleration. Additionally, the effective EoS of DM, $w_{\rm dm}^{\rm eff}$ as displayed in the upper left graph of Fig. 2, becomes negative and approaches 1-1. That means, it behaves like a DE fluid.
@@
For $w=0$, besides the three isolated critical points $A_{2},~A_{3}$ and $A_{4}$ given in Table 1, we find one critical line $A_{7}(x_{c},0)$ for which the critical points $A_{0}$, $A_{1}$ lie in this critical line. In the phase space $R$, one can obtain $z^{\prime}>0$ and $z^{\prime}<0$ for $z<\frac{3}{3+\alpha x}$ and for $z>\frac{3}{3+\alpha x}$ respectively, and $x^{\prime}$ takes always positive value (since $\alpha>0$). By the above argument, one can conclude that the critical line $A_{7}$ shows unstable nature, $A_{1},~A_{2}$ behave like saddle points and $A_{3}$ is the only globally stable point. Here, $A_{3},~A_{4}$ correspond to accelerating phase and $A_{7}$ gives decelerating phase, while decelerating parameter is undefined at $A_{2}$. The phase space diagram is highlighted in the upper middle plot of the Fig. 1 where one can notice that the universe exits from its past decelerating phase and enters into the DM dominated accelerating phase, as depicted in the upper right plot in Fig. 2. We also notice that $w_{\rm dm}^{\rm eff}$ becomes negative and approaches 1-1, as noticed in the earlier case with $w>0$. Along with these results, a set of orbits traverses the decelerating phase $(q>0)$, then enters into an accelerating phase with 1<q<0-1<q<0 but successively these orbits enter into a super accelerating phase (q<1)(q<-1), and finally the acceleration slows down and it continues with q=1q=-1.
@@
When $\frac{1}{3}\leq w<0$, in this interval of $w$, the dynamical system (23) gives five critical points, namely, $A_{0},~A_{1},~A_{2},~A_{3}$ and $A_{4}$. Now we can conclude about the stability of the critical points via investigation of the sign of the eigenvalues and the direction of the flow on the boundary of the unit square. It is clear that $A_{0},~A_{2},~A_{3}$ are saddle by behavior and $A_{1}$ is unstable. Here, $A_{4}$ is the only globally stable point in the phase space. The critical points $A_{3},~A_{4}$ correspond to matter dominated accelerating phases, while $A_{1}$ represents the matter dominated decelerating phase. Note that, $A_{0},~A_{2}$ both are second fluid dominated critical points but again $A_{0}$ shows deceleration and we can not say about the acceleration at $A_{2}$. The phase space trajectories are similar to the upper right plot of the Fig. 1. In this case, the universe quits its past DM dominated decelerating phase and closes in accelerating DM dominated phase. Again, some special orbits tracing accelerating phase (1<q<0)(-1<q<0) and super accelerating phase (q<1)(q<-1), finish in the phase where q=1q=-1. The slowing down of cosmic acceleration is also realized in this case.
@@
When $1<w<-\frac{1}{3}$, this case is similar to the above case where $w$ lies in the interval $\left[-\frac{1}{3},0\right)$. The only difference with the above case is in the decelerating parameter where in the above case we get $q\geq 0$ at $A_{0}$ but here, $q<0$ at $A_{0}$ which represents acceleration. The upper right plot of Fig. 1 exhibits the qualitative features of the critical points.
@@
When w=1w=-1, in this case, we have two critical lines $A_{5}(0,z_{c})$ and $A_{6}\left(x_{c},\frac{3}{3+\alpha}\right)$ along with the critical points $A_{0},~A_{1},~A_{2},~A_{3},~A_{4}$ depicted in the Table 1. Here, $A_{0},~A_{2}$ belong to the critical line $A_{5}$ and $A_{6}$ contains the point $A_{4}$. In the phase space above the $z=3/(3+\alpha)$ line, $x^{\prime}$, $z^{\prime}$ are positive and negative respectively, and below the line $z=3/(3+\alpha)$, $x$ and $z$ are decreasing and increasing respectively. Again, the separatrix $z=\frac{3}{\alpha}(1-z)\sqrt{1-x}$ connecting the critical points $A_{1}$ and $\left(0,3/(3+\alpha)\right)$ divide the phase space below the line $z=3/(3+\alpha)$ into two regions: the orbits below the separatrix converge to the part of the critical line $A_{5}$ where $z<\frac{3}{3+\alpha}$ and the orbits above the separatrix approach the critical line $A_{6}$. As a result, $A_{6}$ behaves like an accelerating late-time attractor where DM and DE both may co-exist. So, the critical point $A_{6}$ has the potentiality to alleviate the coincidence problem. Again, the critical line $A_{5}$ where $z>\frac{3}{3+\alpha}$ are unstable in nature but the part of the critical line $A_{5}$ with $z<\frac{3}{3+\alpha}$ represent late time accelerating attractor dominated by the DE only. The point $A_{3}$ is saddle which shows matter dominated accelerated phase, while $A_{1}$ corresponds to matter dominated decelerating unstable point. The qualitative behavior are highlighted in the lower left plot of Fig. 1. Here, we have two possibilities: one is, above the separatrix, our universe shows the transition from its past DM dominated decelerating phase to the accelerating scaling solutions and on the other hand, below the separatrix, the universe follows DM dominated decelerating phase to completely DE dominated accelerating phase. The lower left plot in Fig. 2 highlights the dynamic behavior of the DM density parameter $\Omega_{\rm dm}^{\rm eff}$, DE density parameter $\Omega_{f}$, decelerating parameter $q$ and $w_{\rm dm}^{\rm eff}$ as functions of $e$-folding number $N=ln(a)$, reflecting key features of the model’s expansion history. One can notice that the effective EoS of DM being slightly negative during the late-evolution of the universe stays close to zero.